• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 18:24:15 on February 26, 2009

    Entry number 263538

    This entry is a followup to: 263534

    keyword=BB Cer PMT 18 TDC cable wasn't seated well -- fixed

    BB Cer PMT 18 cable running to the TDC wasn't seated well at the 50 ohm splitter -- fixed.

    Signals on PMT 12 look good all the way up to the discriminator. There is actually a little background that the y-axis auto-scaling is hiding on the plot so there is actually some background present. We'll have to see how it performs with beam.